# Towards scalable robotic intervention of children with Autism Spectrum Disorder using LLMs

### v1

@@ Abstract
In this paper, we propose a social robot capable of verbally interacting with children with Autism Spectrum Disorder (ASD). This communication is meant to teach perspective-taking using text generated using a Large Language Model (LLM) pipeline. The social robot NAO acts as a stimulator (verbally describes a social situation and asks a question), prompter (presents three options to choose from), and reinforcer (praises when the answer is correct). For the role of the stimulator, the social situation, questions, and options are generated using our LLM pipeline. We compare two approaches: GPT-2 + BART and GPT-2 + GPT-2, where the first GPT-2 common between the pipelines is used for unsupervised social situation generation. We use the SOCIALIQA dataset to fine-tune all of our LLM pipelines. We found that the GPT-2 + BART pipeline had a better BERTscore for generating the questions and the options by combining their individual loss functions. This observation was also consistent with the human evaluations. Lastly, the unsupervised generation of social situations was visualized using T-SNE plots, and the entire pipeline was evaluated for appropriateness for children with ASD by human experts.
Index Terms:
Social HRI, AI-Enabled Robotics, Autonomous AgentsI Introduction
@@
Social robots have had an immense impact on the lives of people in the past decade [1]. They currently share various scenarios of coexistence with humans. This coexistence accounts for Human-Robot Interaction (HRI) in domestic environments like the example of the Roomba robot [2] or of a more interactive robot like Autom that was built to make behavior changes in a more naturalistic user setting [3].
Within healthcare, social robots have undertaken diverse roles too. These roles have required the robot to be involved in a multi-modal interaction with humans. These include engaging in multi-modal interactions with humans, as exemplified by the telepresence robot discussed in [4], facilitating both visual and verbal communication. Other examples of robots being used in healthcare include robots used in surgery [5, 6]. Moreover, robots have found application in psychological interventions, representing another domain where their presence can be immensely notable, as has been seen in ASD interventions [7]. [8].
@@
Autism Spectrum Disorder, as described by the DSM-5, describes individuals that exhibit: 1) challenges in social communication and interaction and 2) restricted and repetitive behaviors [9]. Since there is no cure for ASD, there are countless interventions and treatments, as has been highlighted by the authors in [10]. As robotic intervention of ASD is getting popular with evidences of positive outcomes [11], the development and evaluation of such platforms have gained momentum. Since the population of people with ASD is 1.5%, with every 1 of 54 children in the US alone, sustainability in robotic intervention becomes essential.
@@
Since children with ASD lack an understanding of social cues and situations, a robot that can teach such skills has been a widely proposed solution [12]. So far, the solutions in the literature for robotic intervention have had limited scope for scalability because of the number of finite responses, actions, or pre-defined contents [13] that the robot can generate. Hence, there is a need to bridge this gap with the natural language processing literature that can be used to generate text and speech which the robot can deliver specifically for children with ASD to help them understand various social skills.
To the best of our knowledge, this is the first work in the literature that attempts to increase the scalability of robotic intervention solutions by generating the contents autonomously to teach perspective-taking to children with ASD in a therapy session. In this paper, we make an effort to bridge this gap between language models and their embodiment as a robot for an intervention of children with ASD. In addition, the robot also takes up the role of a prompter and reinforcer (described in Section IV).
@@
This paper has been arranged in the following way: Section II outlines the existing work in the literature related to robots in ASD intervention (Section II-A), robots to teach perspective-taking (II-B), and text generation (Section II-C). Section III highlights the primary contribution of the paper followed by Section IV which describes the workflow highlighting how the robot acts as a stimulator, prompter and reinforcer. Further, Section V describes the problem formulation that we use in this paper. Section VI outlines the results and discussions culminating in Section VII describing the conclusion.
@@ II-A Robots in ASD intervention
The robotic intervention of children with Autism Spectrum Disorder has witnessed variety in both the choice of robot and the respective therapeutic or educational setting. This approach has facilitated the integration of classical methodologies combined with a socially assistive robot, which has urged researchers to examine its effectiveness. In this direction, the authors in [14] have evaluated a Wizard-of-Oz style therapy following the Applied Behavior Analysis (ABA) structure. Based on the efficacy metrics proposed (skill execution and prompt dependency), they could make better conclusions on the clinical merits of a robot-assisted intervention of ASD, which contradicts the HRI metrics that considers gaze, communication, and affect. They also demonstrated the positive effects of such an intervention on one of their study participants. Similar studies done by the authors in [15] have also contributed to positive conclusions about social robots.
@@
However, there are limitations to proving an absolute advantage of using robot-assisted intervention compared to a human-lead session [22]. Although both of these sessions have proven beneficial to fetch positive outcomes, robots, compared to humans, have yet to be considered serious agents owing to the lack of sophistication they can offer [23, 22]. Some of these limitations can be attributed to the type of robot employed in a therapy scenario [23]. Since different kinds of robots have been used so far in various intervention scenarios [21, 24, 25], there is no robot-specific intervention protocol mentioned in the Evidence Based Practices for the ‘Technology-Aided Instruction and Intervention (TAII)’ [11]. Although in [26], a more robot-specific scenario has been described by the authors using the KASPER robot for underlining the roles it can take up in a therapy session, it still possesses limitations in the finite reaction possibilities the robot can generate. The robot’s finite response generation ability also adds to the effort a therapist has to put in during a therapy session. As pointed out in [27], the therapist’s involvement in generating content or controlling the robot, as in a Wizard-of-Oz approach, is not sustainable in the long run. Hence, a more sustainable autonomous solution is needed in which the robot can take up the role of a stimulator that can generate and deliver content during a therapy session [26, 27].
@@ II-C Content generation using Language Models for ASD intervention
Recently, LLMs have been used for robotic specific tasks [48, 49, 50, 51]. This also involves HRI scenarios where goal-oriented LLM usage has been embodied as a robot. Examples include robots being used as a teaching assistant [52], involved in task planning [53] or, being used as an intervention tool in healthcare [54].
@@
Further, generating content using LLMs for a robotic intervention of ASD is still an active area of research. To the best of our knowledge, the only current work that combines a social robot with using LLMs for ASD intervention is [54]. In this, the authors have used a Pepper robot integrating ChatGPT to meet the therapeutic goals of individuals with ASD. However, the structured component used for the problem-solving session is scripted, and the LLM integration is essentially an add-on to provide additional feedback and insights. Hence, it still does not fully meet the autonomy requirements for content generation during therapy sessions, which the authors pointed out in [26].
@@
Hence, the motivation of this paper comes from the absence of a system that can be used for real-time content generation and take another step towards delivering high-end therapeutic scenarios towards a goal-oriented therapy.
III Primary Contributions
The main contributions of this paper are:
To the best of our knowledge, this is the first work that generates content to teach perspective-taking to children with ASD using LLMs.
we compare a sequence-to-sequence model (BART) with a decoder-based model (GPT-2) for text-generation tasks using BERTscore and also verify the outcome using human evaluations.
we evaluate the text generation for appropriateness for children with ASD through human evaluators from the clinical and special education domains.
IV Methodology
In this paper, the robot takes up three roles out of six possible suggested by the authors in [26] with the KASPER robot. Here, instead of the KASPER robot, we have used the NAO robot to take up the role of a provoker, reinforcer, and prompter. The workflow has been shown in Figure 2.
IV-A Role of the robot
To take up the role of the stimulator, the NAO robot generates contexts that are based on perspective-taking scenarios for the child with ASD. Then, NAO generates questions for the child to answer based on those contexts. If the child stays quiet for over 10 seconds, NAO repeats the question. If there is no reply from the child’s end, the robot gives prompts in the form of three options. This prompting also occurs when the child answers incorrectly. However, if the child can answer correctly, the robot praises the child for performing well and then moves on to a new example. This cycle continues until the termination step.
IV-B Role of the therapist
The decision of selecting the accuracy of the answers in both scenarios (one where prompts are given in the form of options and the other when no prompts are given) lies in the hands of the human instructor. The instructor is also responsible for initiating the robot, whether the answers given by the child are correct or not. Only then will the robot switch to the next pipeline segment shown in Figure 2.
@@ IV-B Role of the therapist
A combination of the GPT and BART architectures generates these contexts, questions, and options. For this, we use the SOCIALIQA dataset, which is a social commonsense reasoning benchmark [55]. This dataset has 38,000 data points containing contexts, questions based on those contexts, and multiple-choice options for those questions. The questions broadly fall into three categories. Based on the context, the questions can inquire 1) about motivation, 2) about next steps, or 3) emotional reactions.
V Problem Formulation
@@
where is the generated context by fine-tuned GPT-2. For fine-tuning the BART model to generate questions and the possible answer options, BartForConditionalGeneration is used. Figure 3 shows the model that has been used in this paper. For every context vector , the BART model generates text in the following way:
where is the index of the data points, , , and, are the total number of tokens for the question, option A, option B and, option C, respectively. Further, , , and, are the predicted question, option A, option B and option C respectively. As can be seen from Figure 3, the same BART model is fine-tuned with a combination of different training inputs. Equations 2a-2d give a hierarchical nature to the training where the context is first used to predict the question and so on. For each of these predictions, the cross entropy loss is calculated and added for backpropagation:
@@
This approach was also compared to the unsupervised GPT-2 generation of the question and three options based on the context. We introduce three special tokens for controlled text generation using GPT-2. In addition to the and tokens, we also use <context>:, <question>:, <ansa>:, <ansb>: and, <ansc>: to control the generation process of each of the individual parts for each data point. Once these tokens have been added to the tokenizer and the dataset tokenized, we then fine-tune the GPT-2 model and use the cross entropy loss. Each data point before tokenization can be represented as:
VI Results and Discussions
As discussed previously, we have used the SOCIALIQA dataset for all of our generation tasks. We used the premium GPU from Google Colab with A100 GPUs for training. The dataset was split into train (75%), eval (15%), and test data (10%). Adam was chosen to be the optimizer. The SOCIALIQA dataset has one context, one question, and three options for every data point. So, we fine-tuned GPT-2 for generating these contexts and then fed the output of the GPT-2 into our BART pipeline explained in the equation V to generate the questions and the options. We compared this pipeline with a similar pipeline, replacing BART with GPT-2. Further, we also present human evaluations later to validate the results and the choice of the pipeline.
VI-A Context generation using GPT-2
For generating the context, we fine-tuned a GPT-2 model based on the contexts in the SOCIALIQA dataset. To demonstrate how similar the generated contexts were to the test set, we took one thousand GPT-2 generated texts (after fine-tuning) and the same number of data points from the test set. Further, we used the T-SNE plots to visualize how the test subset and the GPT-2 generated set overlap. We vectorized the generated texts and the subset of the test dataset using scikit-learn’s TfidfVectorizer to convert both the test and the generated datasets into a tf-idf feature matrix [56]. Once vectorized, the T-SNE plot was used to visualize the high-dimensional text data into two components. Since we wanted to compare how similar the GPT-2 generated texts and the actual test set, we just used the number of components to differentiate between them. As seen from Figure 4a, the generated and the test set have a very large overlap. To characterize this overlap, we considered the data points of the scatter plot and used a Support Vector Classifier (SVC) with a radial basis kernel to see if the classifier could separate the data points into two distinct classes. Figure 4b and 4c show that the classification accuracy is 0.51, and the F1-score is around 0.55, too. This low F1-score signifies that the classifier could not very distinctly classify the values of the T-SNE plot (essentially the low-dimensional embeddings of the high-dimensional text data) [56]. The low scores on the classifying metrics signify that the text generated by fine-tuning GPT-2 generates data similar to the test set.
VI-B Evaluation metrics
We use BERTscore to evaluate our generation task [57]. We do this across generations of questions, options, and both of them combined. Other metrics are used in the literature, like the BLEU score [58]. The reason for using BERTscore is because it uses cosine similarity instead of just making an n-gram-based comparison between the actual text and the generated ones. In addition, for calculating the cosine similarities, it uses BERT embeddings between each generated token and every reference token [57]. This ensures not only exact phrase matching between the reference and the candidate text, but also the contextual relevance between the two. Since there are different BERT-based models to calculate the BERTscore from, we choose three of those models: 1) ‘distillery-base-uncased’, 2) ‘roberta-base’ and 3) ‘microsoft/deberta-large-mnlimicrosoft/deberta-xlarge-mnli’. Among these, microsoft/deberta-xlarge-mnli is currently the closest to human evaluations according to [59].

### v3

# Human-mediated Large Language Models for Robotic Intervention in Children with Autism Spectrum Disorders

@@ Abstract
The robotic intervention for individuals with Autism Spectrum Disorder (ASD) has generally used pre-defined scripts to deliver verbal content during one-to-one therapy sessions. This practice restricts the use of robots to limited, pre-mediated instructional curricula. In this paper, we increase robot autonomy in one such robotic intervention for children with ASD by implementing perspective-taking teaching. Our approach uses large language models (LLM) to generate verbal content as texts and then deliver it to the child via robotic speech. In the proposed pipeline, we teach perspective-taking through which our robot takes up three roles: initiator, prompter, and reinforcer. We adopted the GPT-2 BART pipelines to generate social situations, ask questions (as initiator), and give options (as prompter) when required. The robot encourages the child by giving positive reinforcement for correct answers (as a reinforcer). In addition to our technical contribution, we conducted ten-minute sessions with domain experts simulating an actual perspective teaching session, with the researcher acting as a child participant. These sessions validated our robotic intervention pipeline through surveys, including those from NASA TLX and GodSpeed. We used BERTScore to compare our GPT-2 BART pipeline with an all GPT-2 and found the performance of the former to be better. Based on the responses by the domain experts, the robot session demonstrated higher performance with no additional increase in mental or physical demand, temporal demand, effort, or frustration compared to a no-robot session. We also concluded that the domain experts perceived the robot as ideally safe, likable, and reliable.
I Introduction
@@
Social robots have positively impacted people’s lives in the past decade [1]. They currently share various scenarios of coexistence with humans. This coexistence accounts for Human-Robot Interaction (HRI) in domestic environments like the Roomba robot [2], or of a more interactive robot like Autom that was built to make behaviour changes in a more naturalistic user setting [3].
@@
Within healthcare, social robots have undertaken diverse roles too. These roles have required the robot to be involved in a multi-modal interaction with humans. Examples include the telepresence robot discussed in [4], facilitating both visual and verbal communication. Other examples of robots being used in healthcare include robots used in surgery [5, 6]. Moreover, robots have found application in psychological interventions, representing another domain where their presence can be immensely notable, as has been seen in ASD interventions [7]. [8]. Autism Spectrum Disorder, as described by the DSM-5, describes individuals that exhibit: 1) challenges in social communication and interaction and 2) restricted and repetitive behaviors [9]. Since there is no cure for ASD, there are countless interventions and treatments, as has been highlighted by the authors in [10]. As robotic intervention of ASD is getting popular with evidence of positive outcomes [11], the development and evaluation of such platforms have gained momentum. Since the population of people with ASD is 1.5% [12], with every 1 of 54 children in the US alone, sustainability in robotic intervention becomes essential.
@@
Since children with ASD cannot understand social cues and situations, a robot that can teach such skills has been a widely proposed solution for both verbal and non-verbal communications [13, 14]. So far, the solutions in the literature for robotic intervention have had limited scope for scalability because of the number of finite responses, actions, or pre-defined contents [15] that the robot can generate. Hence, there is a need to bridge this gap between the deep learning literature that can be used to generate content (speech, motions, etc.) that the robot can deliver specifically for children with ASD to help them understand various social skills.
To the best of our knowledge, this work is novel in its approach to creating an embodied speech AI agent that can independently generate verbal content during a perspective-taking teaching session for children with ASD. This introduction of autonomy in classical human-led interventions or even teleoperated robot-led interventions shows the potential for more general interventions aimed at individuals with ASD. We compared our GPT-2 BART pipeline to an all GPT-2 pipeline and found the former to be better for question, option, and overall generation of both questions and options (explained more in Section V). In addition to this technical contribution, we also conducted ten-minute sessions with domain experts and a child actor, with the robot acting as an initiator, prompter, and reinforcer. Based on the survey data collected at the end of each of these ten-minute sessions, we found that the robot session demonstrated higher performance with no additional increase in mental demand, physical demand, temporal demand, effort, or frustration as compared to a no-robot session (NASA TLX data). We also concluded that the domain experts perceived the robot as ideally safe, likable (GodSpeed data), and the entire session to be technically relevant and reliable (Appropriateness).
@@
This paper has been arranged in the following way: Section II outlines the existing work in the literature related to robots in ASD intervention (Section II-A), robots to teach perspective-taking (II-B), and text generation (Section II-C). Section III which describes the workflow highlighting how the robot acts as an initiator, prompter and reinforcer. Further, Section IV describes the problem formulation that we use in this paper. Section V outlines the results and discussions, followed by Section VI, which describes the limitations and scope for future work. Finally, Section VII describes the conclusion.
@@ II-A Robots in ASD intervention
The robotic intervention of children with ASD has witnessed variety in both the choice of robot and the respective therapeutic or educational setting. This approach has facilitated the integration of classical methodologies combined with a socially assistive robot, which has motivated researchers to examine its effectiveness. In this direction, the authors in [16] have evaluated a Wizard-of-Oz style therapy following the Applied Behavior Analysis (ABA) structure. Based on the efficacy metrics proposed (skill execution and prompt dependency), they could make better conclusions on the clinical merits of a robot-assisted intervention of ASD, which contradicts the HRI metrics that considers gaze, communication, and affect. They also demonstrated the positive effects of such an intervention on one of their study participants. Similar studies done by the authors in [17] have also contributed to positive conclusions about social robots.
@@
Although both the robot-led and human-led sessions have proven beneficial to fetch positive outcomes, robots, compared to humans, have yet to be considered serious agents owing to the lack of sophistication they can offer [24, 25]. Although in [26], the authors used the KASPER robot to underline the roles it can take up in a therapy session, it still has limitations in the finite reaction possibilities the robot can generate. The robot’s finite response generation ability also adds to the effort a therapist has to put in during a therapy session. As pointed out in [27], the therapist’s involvement in generating content or controlling the robot, as in a Wizard-of-Oz approach, is not sustainable in the long run. Hence, a more sustainable autonomous/semi-autonomous solution is needed in which the robot can take up the role of an initiator that can generate and deliver content during a therapy session [26, 27].
@@ II-C Content generation using Language Models for ASD intervention
Recently, LLMs have been used for robotic-specific tasks [48, 49, 50]. This also involves HRI scenarios where goal-oriented LLM usage has been embodied as a robot. Examples include robots being used as a teaching assistant [51], involved in task planning [52] or, being used as an intervention tool in healthcare [53].
@@
Further, generating content using LLMs for a robotic intervention of ASD is still an active area of research. To the best of our knowledge, the only work that combines a social robot with using LLMs for ASD intervention is [53]. In this, the authors have used a Pepper robot integrating ChatGPT to meet the therapeutic goals of individuals with ASD. However, the structured component used for the problem-solving session is scripted, and the LLM integration is essentially an add-on to provide additional feedback and insights. Hence, it still does not fully meet the autonomy requirements for content generation during therapy sessions, which the authors pointed out in [26].
@@
Hence, this paper is motivated by the absence of a system that can be used for real-time content generation and take another step towards delivering high-end therapeutic scenarios in a goal-oriented therapy.
III Methodology
III-A Human mediated autonomy
In the pipeline shown in Figure 3, the robot takes up three roles: initiator, prompter, and reinforcer similar to the definitions suggested by the authors in [26]. The session starts with user input 1, where the domain expert is asked through a graphical user interface (GUI) on the computer if they want to continue the session. If they select yes, they are then shown the context (which describes a social situation), question and the respective options. If they feel that this set of context, question and, options are appropriate for children with ASD, then they press yes. This allows the robot to take the role of the initiator, where it speaks the context and the question only and waits for a response from the child.
This generation of context and the question is not pre-determined. Rather, it is generated by our LLM model (described in section IV). If the child answers the question correctly (decided by the user input 3), the robot acts as a reinforcer and provides positive reinforcement by praising them on a job well done. However, if the child does not answer or answers incorrectly (decided by the user input 3), the robot takes up the role of the prompter.
In the role of a prompter, the robot repeats the context and the question and now speaks out the options to choose from. If the child answers correctly (decided by the user input 4), the robot gives positive reinforcement at this stage by congratulating them on correctly answering. On the other hand, if the user input 4 is either ‘no’ or ‘no response’, the robot returns to user input 1 and repeats this cycle.
As described above, this is a semi-autonomous system, meaning the autonomy lies in the content generation part. However, the decisions on whether the content is appropriate for children with ASD and whether the child answers correctly are made by the domain expert through a computer. This establishes control over what the robot can do, which is essential given a population such as children with ASD.
III-B Autonomous content generation using LLMs
As shown in Figure 5, we use two LLM pipelines to generate the questions and the options. However, the context generation through fine-tuned GPT-2 remains the same for both pipelines. The problem formulation is explained in Section IV.
@@ III-B Autonomous content generation using LLMs
For fine-tuning both of our LLM pipelines, we use the SOCIALIQA dataset, which is a social commonsense reasoning benchmark [54]. This dataset has 38,000 data points containing contexts, questions based on those contexts, and multiple-choice options for those questions. The questions broadly fall into three categories. Based on the context, the questions can inquire 1) about motivation, 2) about next steps, or 3) emotional reactions.
III-C Data collection
In this study, we not only make a technical contribution by leveraging LLMs to introduce autonomy in perspective-taking sessions using a robot but also collect the opinions of domain experts on some surveys. We asked four domain experts to perform a 10-minute session each (independently) with the robot and a participant (not an actual child with ASD but a child actor). The domain experts included one psychologist, one neurophysician, one professor from the special education domain, and a clinical assistant professor from the special education domain, each having substantial experience conducting sessions in therapeutic, clinical, and educational settings with children with ASD. At the end of the 10-minute session, we asked them to fill out the NASA Task Load Index (NASA TLX), comparing their session with the robot with classical non-robotic sessions for perspective-taking. In addition, we asked them to fill out the Godspeed survey to understand their perceptions of the robot. Lastly, we prepared an additional survey (called the Appropriateness survey) specific to our session to understand how useful our human-mediated autonomous robotic pipeline was for teaching perspective-taking to children with ASD.
IV Problem Formulation
@@
where is the generated context by fine-tuned GPT-2. For fine-tuning the BART model to generate questions and the possible answer options, BartForConditionalGeneration is used. Figure 6 shows the model that has been used in this paper. For every context , the BART model generates text in the following way:
where is the index of the data points, , , and, are the total number of tokens for the question, option A, option B and, option C, respectively. Further, , , and, are the predicted question, option A, option B and option C respectively. As can be seen from Figure 6, the same BART model is fine-tuned with a combination of different training inputs. Equations 2-2 give a hierarchical nature to the training where the context is first used to predict the question and so on. For each of these predictions, the cross entropy loss is calculated and added for backpropagation:
@@
This approach was also compared to the GPT-2 generation of the question and three options based on the context. We introduce three special tokens for controlled text generation using GPT-2. In addition to the and tokens, we also use <context>:, <question>:, <ansa>:, <ansb>: and, <ansc>: to control the generation process of each of the individual parts for each data point. Once these tokens have been added to the tokenizer and the dataset tokenized, we then fine-tune the GPT-2 model and use the cross entropy loss. Each data point before tokenization can be represented as:
V Results and Discussions